OpenStudy (sarahc):

Which of the following atoms has the same number of neutrons as 19F? 16O 18F 20Ne 23Na

OpenStudy (sarahc):

Im thinking c?

OpenStudy (anonymous):

19 F has 19 - 9 = 10 neutrons 20 Ne = 20-10 = 10

OpenStudy (anonymous):

correct its C

OpenStudy (sarahc):

thank you!

OpenStudy (anonymous):

post this in chemistry, not math.

OpenStudy (anonymous):

19 F has 19 - 9 = 10 neutrons O: 16-8 =8 Ne = 20-10 = 10 ANSWER: Neon!!

OpenStudy (sarahc):

I have another :) An element has three naturally occurring isotopes. Use the information below to calculate the weighted average atomic mass of the element, showing both the setup and the final answer for the calculation. Isotope Atomic Mass Percent Abundance X 1.01 u 99.984% Y 2.01 u 0.014% Z 3.02 u 0.002%

OpenStudy (sarahc):

@surdawi

OpenStudy (anonymous):

average atomic mass of the element= 1.01(99.984/100) + 2.01(0.014/100) + 3.02(0.002/100) = 1.01 u